OpenStudy (anonymous):

f(x)=3x and g(x)=-2x-6 find (fog)(4)

OpenStudy (kmeis002):

Do you know how to find \( f \circ g \) first?

OpenStudy (anonymous):

no i dont really even know what that means tbh

OpenStudy (isaidavila):

F(x)=3x G(x)=-2x-6 So (F°g) will be 3(-2x-6) -6x-18 So find for x=4 -6(4)-18 -24-18=??

OpenStudy (anonymous):

thank you so much

OpenStudy (kmeis002):

The notation \( f \circ g\) represents functional composition, meaning you take the output of \(g\) and use that as the input of \(f\). So \[ f \circ g(x) = f(g(x)) = 3(-2x-6) = -6x-18\]

OpenStudy (kmeis002):

Then once you have that, you simply plug in 4.

OpenStudy (anonymous):

thank you